# Lost in Space: Optimizing Tokens for Grammar-Constrained Decoding

### v1

@@ 5 Analysis
Previously we presented the correlation between model outputs constrained by various output formats and human labels. Figure 3 shows the correlations between these formatted outputs themselves for two model families, Llama and Phi. We collect observations produced with each token family twice: once for numerical values, and again for word values. Stable intra-token set correlations should yield a solid color in a correlation matrix. We instead find all correlation matrices exhibit poor intra-token set regularity. This indicates no model in our study has achieved stable token representations for the tasks and token sets tested.
Neither do output formats belonging to the same token family (from a human perspective) have correlated outputs, again indicating the models tested have not achieved token set stability. Comparing small and large models reveals increasing parameter count (and consequently training tokens) shows less differentiation between output formats for Llama, Gemma, and Qwen when increasing scale. This does not improve consistency for Phi 3. One possible explanation for this phenomenon is the large portion of synthetic data used in the Phi training data; synthetic data can amplify certain signals in noise, increasing the possibility of overfitting during pre-training.
Whitespace in token embeddings.
We find that formats that start with a space have better zero-shot performance than otherwise identical formats that do not have that space. To explain this difference we ask, to what degree do tokens incorporating LW (henceforth LW tokens) differ from their non-LW twins in the embedding space? We obtain the tokenizers for each of our model families and extract all LW/non-LW token pairs from their corresponding vocabularies. All vocabularies contain token pairs. We find between $22\%$ and $32\%$ of tokens in the Phi, Llama and Qwen vocabularies participate in a pair, while $40\%$ of the Gemma vocabulary participates in a pair. Put another way, the lower bound observed has one in ten tokens having a twin — and the upper bound is every one in $2.5$. Do these pairs share semantic representations? To test this, we take the token index keys for each pair and obtain the corresponding columns of the Llama 3.1 8b and 70b embedding matrices. Increasing model size slightly increases pair similarity. The cosine similarity of pairs is $0.369$ for 8b and $0.386$ for 70b, with an effect size of $d\!=\!0.13$. To verify that this effect is not the result of all token embeddings becoming more similar as we increase model size, we calculate cosine similarity between $20,000$ random pairs of token embeddings for 8b and 70b. In fact, the average pair of vectors is less similar in the larger model: pairs sourced from 8b have a mean similarity of $0.016$ versus $0.008$ for 70b, with an effect size of $d\!=\!-0.38$.
A t-SNE projection (Van der Maaten and Hinton, 2008) of LW and non-LW token embeddings is shown in Figure 4 for small and large Llama models. Increasing model scale (and correspondingly training time assuming compute optimality) results in vectors that t-SNE is better able to distribute uniformly. Furthermore, we can observe LW/non-LW tokens co-occupy space to a greater degree as scale increases, suggesting our measured cosine similarities do in fact indicate more exchangeable representations.
Our results indicate LW/non-LW token pairs are not wholly exchangeable on an embedding basis. We turn to the training data to investigate one potential cause for this phenomenon.
@@
Counting token occurrences furthermore reveals LW tokens are consistently more prevalent than their non-LW token duplicates with all tokenizers. We find the Llama, Gemma, and Qwen tokenizers emit LW tokens at rates from $2.5$ to $2.7$ time greater than non-LW tokens, indicating our models of interest were exposed to many more LW tokens than non-LW tokens during pre-training. Phi is again an outlier, with its tokenizer producing LW tokens at a rate $1.66$ times greater than non-LW tokens. Our results provide one possible explanation for why incorporating LW into grammars leads to increased classification performance: the LW tokens were observed at a greater frequency during pre-training. The Phi tokenizer producing fewer LW tokens relative to non-LW tokens supports this conclusion, given we also observe Phi showing limited response to newlines and leading whitespace.
@@ 6 Discussion
Conditioning LLM results on specific output formats is a promising and popular way to ensure valid annotation results, but can reduce performance. We systematically test a range of familiar formats that appear semantically similar to humans, along with a set of format options that users may not even be aware of. We find there are substantial differences between seemingly identical formats. By varying over tokens differentiated by syntactic – but not semantic – qualities, we show GCD-guided LLMs are sensitive to features like leading whitespace, numerical representation, and newlines. By better evaluating and surfacing these differences, we hope to close the gap in output-filtered performance without modifying existing decoding systems.
We recommend using as many of the following strategies as possible:
@@
Prefer numeric outputs.
We recommend selecting token ranges matching the intended output format. For example, if the classification task at hand requires numerical values be selected, then it is important the corresponding grammar incorporate numerical tokens. Subword token representations in existing LLMs are not yet coherent enough to dependably exchange numerical tokens for their semantically-equivalent word tokens, or vice versa. While future models may one day reach representation parity, it is recommended to select context-appropriate patterns until then.
@@
It is helpful to offer the model the greatest degree of choice if possible. Consider a grammar guiding models to emit a scalar value – our results suggest it is beneficial to offer the model a choice from the closed interval $[0,100]$ rather than $[1,10]$. It is further helpful to represent this range with the fewest tokens possible to prevent errors propagating over multiple inferences.
@@
Finally, larger models, when available, are usually better. They are generally — though not always — more capable and are typically less influenced by differences in requested output formats.
@@ 7 Conclusion
Grammar-constrained decoding (GCD) has the potential to close the gap between the task-specificity of fine-tuned models and the rapid iteration of zero-shot models. But users must not only write good prompts, they must also write effective output-grammar constraints. This work indicates that users may unintentionally knock models out of their “comfort zone” by inadvertently selecting patterns that are misaligned with pre-training or RLHF fine-tuning. We find that better understanding the effect of subtle format differences has significant benefits and opens new opportunities for research.
Advances in quantization, software infrastructure, and model training have made powerful LLMs increasingly available on laptop-grade hardware. Of the eight LLMs studied in this work, six will load on computer systems containing 32GB of RAM. The remaining two, Llama 70b and Qwen 72b, require a minimum 48GB assuming quantization is kept to a minimum of six bits per weight. We find smaller models are more liable to produce non-valid output, meaning GCD is especially valuable for researchers working in memory-constrained environments. We also find that smaller models show more sensitivity to GCD formats. The results of this work can help to ensure that the benefits of low-compute LLMs are fully realized.
The increasing popularity of structured output indicates there is appetite for employing LLMs as pseudo-classifiers. GCD is critical for ensuring the success of structured output, and we see our current brute-force method as a promising initial step. Generating $1.28$ million generations on dual RTX A6000 is a time-consuming process. Future work should consider how else GCD optimality can be determined. One interesting pathway will be to further study how subword token representations impact downstream performance. Can modifications to the pre-training process ensure more efficient transfer of representation between visually identical tokens?

### v2

@@ 5 Analysis
How stable are our models’ responses across format types? We demonstrate the correlations between all annotations made with each format type for two model families, Llama and Phi, in Figure 2. Stable responses with each format should yield a solid color in our correlation matrix — but we instead find all model families (especially their smaller siblings) demonstrate poor intra-format regularity. Neither when we represent the same scales (e.g. integers) both numerically and in letters (e.g. the number 5 versus “five”) do we find correlated annotations. This result indicates no model in our study has achieved stable token representations for the tasks and token sets tested. Comparing small and large models does reveal increasing parameter count (and thereby training tokens) decreases differentiation between formats for all pairs considered belonging to Llama, Gemma, and Qwen. One possible explanation for Phi being an outlier is the large portion of synthetic data used in its training data — synthetic data can amplify certain signals in noise, increasing the possibility of overfitting.
Downstream impact.
Do our results hold for downstream tasks, such as classification and reasoning? If our results in section 4 hold, then we would expect Llama to continue exhibiting non-uniform performance when guided to emit different formats. To investigate, we ran Llama 3.1 8b on seven common LLM reasoning benchmarks with similar formats and treatments as previously described in section 3. We test the following formats: Stock uses letter choices (A, B…), As Integer uses numbers (1, 2…), As Real uses decimal numbers (0.00, 0.33…), As Word uses choice labels (“Choice 1”, “Choice 2”…), and finally With Space and With Newline add leading whitespace and newlines, respectively. We present our findings in Table 5. While all the benchmarks are multiple choice, they vary by the number of alternative answers for each question: BoolQ and Winogrande are binary (Clark et al., 2019; Sakaguchi et al., 2019); ARC-C, MMLU, OpenBookQA, and HellaSwag all present four choices (Clark et al., 2018; Hendrycks et al., 2021; Mihaylov et al., 2018; Zellers et al., 2019); and CommonsenseQA offers five (Talmor et al., 2019). We vary the number of options accordingly. We find Llama 3.1 8b continues to demonstrate uneven performance across formats and treatments, with newlines and leading whitespace improving over stock performance by an average 37% and 10%. Whereas we previously found models preferred real numbers when attempting prediction tasks, here we observe models preferring letters over integers (-22%) and real numbers (-39%) when selecting multiple choice answers. This again suggests models prefer certain formats, with the conventional format of the task (e.g. real numbers for prediction) being a predictor.
Leading whitespace prevalence rates.
@@
[…]
Leading whitespace shared semantics.
Do these pairs share semantic representations? We take the token index keys for each pair and obtain the corresponding columns of the Llama 3.1 8b and 70b embedding matrices. We then take the cosine similarity of each pair. We find increasing model size slightly increases pair similarity. The average cosine similarity of pairs is 0.369 for 8b and 0.386 for 70b, with the distance of the two distributions being $d\!=\!0.13$. To verify that this effect is not the result of all token embeddings becoming more similar as we increase model size, we calculate cosine similarity between 20,000 random pairs of token embeddings for 8b and 70b. In fact, the average pair of vectors is less similar in the larger model: pairs sourced from 8b have a mean similarity of 0.016 versus 0.008 for 70b (Cohen’s $d\!=\!-0.38$).
A t-SNE projection (Van der Maaten and Hinton, 2008) of all LW/non-LW pairs in the embedding matrices of both Llama models is shown in Figure 3. Both projections integrate all tokens not belonging to a pair as a baseline. Increasing model scale (and corresponding training time assuming compute optimality) results in vectors that t-SNE is better able to distribute uniformly. We furthermore observe LW/non-LW tokens co-occupy space to a greater degree as scale increases, which suggests our measured cosine similarities suggest more exchangeable representations as model scale increases. This indicates LW/non-LW token pairs are not wholly exchangeable on an embedding basis.
@@
Counting token occurrences furthermore reveals LW tokens are consistently more prevalent than their non-LW token duplicates with all tokenizers. We find the Llama, Gemma, and Qwen tokenizers emit LW tokens at rates from 2.5 to 2.7 times greater than non-LW tokens, indicating our models of interest were exposed to many more LW tokens than non-LW tokens during pre-training. Phi is again an outlier, with its tokenizer producing LW tokens at a rate 1.66 times greater than non-LW tokens. Our results provide one possible explanation for why incorporating LW into grammars leads to increased classification performance: the LW tokens were observed at a greater frequency during pre-training. The Phi tokenizer producing fewer LW tokens relative to non-LW tokens supports this conclusion given we also observe Phi showing limited response to leading whitespace.
@@ 6 Discussion
Conditioning LLM results on specific output formats is a promising and popular way to ensure valid annotation results, but can reduce performance. We systematically test a range of familiar formats that appear semantically similar to humans, along with a set of format options that users may not even be aware of. We find there are substantial differences between seemingly identical formats. By varying over tokens differentiated by syntactic (but not semantic) qualities, we show GCD-guided LLMs are sensitive to features like leading whitespace, numerical representation, and newlines. By better evaluating and surfacing these differences, we hope to close the gap in output-filtered performance without modifying existing decoding systems.
We recommend the following strategies:
@@
Prefer the conventional format.
We recommend selecting token ranges matching the intended output format. If the task at hand requires numerical values be selected, then it is important the corresponding grammar incorporate numerical tokens. Subword token representations in existing LLMs are not yet coherent enough to dependably exchange numerical tokens for their semantically-equivalent word tokens, or vice versa. It is recommended to select context-appropriate patterns until models reach representation parity.
@@
It is helpful to offer the model the greatest degree of choice if possible. Consider a grammar guiding models to emit a scalar value — our results suggest it is beneficial to offer the model a choice from the range [0,100][0,100][ 0 , 100 ] over [1,5][1,5][ 1 , 5 ]. It is further helpful to represent this range with the fewest tokens possible to prevent errors propagating over multiple inferences.
@@
Finally, our results indicate larger models, when available, are usually better. They are generally (although not always) more capable and are typically less influenced by differences between output formats.
@@ 7 Conclusion
Models have token and formatting preferences that are not predictable nor transparent to users. As trends move towards smaller local models — and increasing GCD-based constraints to compensate — this issue will become more pressing. Going forward, these format preferences should be identified and reported on a model-by-model basis.
Grammar-constrained decoding (GCD) has the potential to close the gap between the task-specificity of fine-tuned models and the rapid iteration of zero-shot models. But users must not only write good prompts, they must also write effective grammars. This work indicates that users may unintentionally knock models out of their “comfort zone” by inadvertently selecting patterns that are misaligned with pre-training or RLHF fine-tuning. We find that understanding the effect of subtle format differences has significant benefits and opens new opportunities for research.
Advances in quantization, software infrastructure, and model training have made powerful LLMs increasingly available on laptop-grade hardware. Of the eight LLMs studied in this work, six will load on computer systems containing 32GB of RAM. The remaining two, Llama 70b and Qwen 72b, require a minimum 48GB assuming quantization is kept to a minimum of six bits per weight. We find smaller models are more likely to produce non-valid output, meaning GCD is especially valuable for researchers working in memory-constrained environments. We also find that smaller models show more sensitivity to GCD formats. The results of this work can help to ensure that the benefits of low-compute LLMs are fully realized.
The increasing popularity of structured output indicates there is appetite for employing LLMs as classifiers. GCD is critical for ensuring the success of structured output, and we see our current brute-force method as a promising initial step in enabling consistent success. Generating 1.281.281.28 million generations is a time-consuming process. Future work should consider more direct methods of assessing GCD optimality. One pathway will be to study how training impacts subword token representations coherence. Could modifying pre-training help ensure more consistent representations for similar tokens?